# Estimating the linear relation between variables that are never jointly observed

### v3

@@ 9 Discussion
This work focuses on a statistical framework designed for extracting dependencies from experiments, specifically introducing linear regression estimators in the context where the predictor and the predicted variables are never jointly observed. In this work we chose the basic linear multivariate setting, prioritizing simplicity and computational feasibility. Particularly, the estimator based on the method of moments makes no hypotheses on data distribution and can be calculated explicitly. The estimator based on optimal transport includes a simple optimization problem, and is based on the Gaussian form of the Wasserstein distance but does not technically require the data to be Gaussian, seeking to approximate them with Gaussian variables in whatever case.
However, these approaches are inapplicable in the cases where linear relationship hypothesis cannot be satisfied. For instance, it is the case with predicting survival data with some continuous biomarker, which is of particular interest in research into radio-induced adverse effects. To be able to consider such scenarios, our model can be extended to a more general case, namely with generalized linear model. The optimal transport estimator appears promising in this context given the fact that the Wasserstein distance allows to compare probability distributions of different nature (for example, continuous and discrete).
Another further research direction lies in investigating alternative methods based on integrated likelihood and Bayesian approaches, which are likely to produce better results in many cases but require putting priors on distributions.
Finally, it would be of interest to work on improving theoretical properties of our estimators for finite samples, namely correcting the negative bias that appears for both estimators. The latter is particularly important in the case of the optimal transport estimator, which may be arising naturally with the Wasserstein distance (Wasserman,, 2023). Correcting this bias would considerably improve the estimator, making it competitive with the approaches mentioned earlier that make numerous assumptions on the data.
Software availability
Python code and the experimental data are freely available at github.com/parsenteva/vivo_lm.
@@ Acknowledgements
The authors gratefully acknowledge the funding from the European Union through the PO FEDER-FSE Bourgogne 2014/2020 programs as part of the project ModBioCan2020, and from the Institut de Radioprotection et de Sûreté nucléaire as part of the project ROSIRIS. The authors also warmly thank Dr. Agnès François and Dr. Annaïg Bertho for the experimental data, as well as Dr. Patrick Tardivel for his numerous remarks that helped to improve the presentation of the manuscript.
Conflict of Interest: none declared.
@@ Appendix A Some classical theorems in asymptotic statistics
Under additional hypotheses, we also get the asymptotic normality of the sequence of estimators $\widehat{\theta}_{n}$ of $\theta_{0}$. we denote by $\nabla_{00}Q_{n}(\theta)$ the Hessian matrix of functional $Q_{n}$ evaluated at $\theta$.
Theorem A.4.
@@
Theorem A.5.
(CLT for bootstrap means)
@@ Appendix B Proofs
The law of large numbers gives that $\widehat{\boldsymbol{\Gamma}}_{X}^{2}\to\boldsymbol{\Gamma}_{X}^{2}$ and $\widehat{\sigma}_{Y}^{2}\to\sigma_{Y}^{2}$ in probability and we deduce, with another application of the continuous mapping theorem, that $\widehat{\sigma}_{Y}^{2}-\widehat{\boldsymbol{\beta}}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{2}\widehat{\boldsymbol{\beta}}\to\sigma_{Y}^{2}-\boldsymbol{\beta}^{\top}\boldsymbol{\Gamma}_{X}^{2}\boldsymbol{\beta}=\sigma_{\epsilon}^{2}$ in probability as $n_{\min}$ tends to infinity. ∎
Proof.
of Lemma 4.2
@@
$\left|\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}\right|$ $\leq C_{2,k}\left\|\widehat{\boldsymbol{\Gamma}}_{X}^{k}\right\|_{sp}\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|.$ (B.1)
@@
$\left|\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right|$ $\leq B_{n}^{k}\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)$ (B.2)
@@
$\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k}+A_{n}^{k}\right),$ (B.3)
@@
We conclude the proof by recalling that $\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ attains its unique minimum at $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ if assumption $\mathbf{H}_{1}$ is fulfilled, so that $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W})\to(\boldsymbol{\beta},\sigma^{2}_{\epsilon})$ in probability in view of Theorem 2.1 in Newey and McFadden, (1994). ∎
Proof.
of Proposition 4.1
@@
$\sqrt{n}\begin{pmatrix}\widehat{\boldsymbol{\mu}}_{X}^{1}-\boldsymbol{\mu}_{X}^{1}\\ \vdots\\ \widehat{\boldsymbol{\mu}}_{X}^{K}-\boldsymbol{\mu}_{X}^{K}\\ \widehat{\mu}_{Y}^{1}-\mu_{Y}^{1}\\ \vdots\\ \widehat{\mu}_{Y}^{K}-\mu_{Y}^{K}\end{pmatrix}$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\mu}\right)$ (B.4)
@@
Application $g$ is differentiable at $\boldsymbol{\theta}=(\boldsymbol{\mu}_{X}^{1},\ldots,\boldsymbol{\mu}_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$, with non null Jacobian matrix denoted by $\mathbf{J}_{\theta}$ (see Chapter 8 and more particularly Theorem 8.3 in Magnus and Neudecker, (2019)). The application of the Delta method (see Theorem 3.1 in van der Vaart, (1998)) permits to get the asymptotic normality convergence result,
@@
where $\boldsymbol{\Gamma}_{\beta_{M}}=\mathbf{J}_{\theta}\boldsymbol{\Gamma}_{\mu}\mathbf{J}_{\theta}^{\top}$. ∎
Proof.
of Proposition 4.2
@@
The proof consists in checking the different points of Theorem A.4. Point (i) is satisfied by the hypotheses, and the point (ii) follows directly from the fact $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ is twice-differentiable in a neighborhood of $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$. To show that (iii) is fulfilled, we consider the following expansion, based on the empirical version of the gradient of $\varphi$,
@@
$\nabla\varphi_{n}$ $=\begin{pmatrix}-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ \sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix}$ (B.5)
@@
Since model (2.1) holds, $\nabla\varphi=0$ and $\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k})-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}^{k}-\boldsymbol{\mu}_{X}^{k}\right)$, we thus deduce with (B.4) the asymptotic normality of the first component of the gradient $\nabla\varphi_{n}$, that is to say $\sqrt{n}\left(-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\right)$ converges in distribution to a centered Gaussian distribution. As far as the second component is concerned, it can be noted that $\widehat{\boldsymbol{\Gamma}}_{X}^{k}$ converges in probability to $\boldsymbol{\Gamma}_{X}^{k}$ and by the continuous mapping theorem, $\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}\to\sigma_{Y,k}$ in probability. It can also be noted that, under the moment condition ${\mathbb{E}}\left[\|\mathbf{X}\|^{4}|G=k\right]<\infty$, the central limit theorem gives that $\sqrt{n}\left(\widehat{\boldsymbol{\Gamma}}_{X}^{k}-\boldsymbol{\Gamma}_{X}^{k}\right)$ converges in distribution to a centered Gaussian multivariate distribution, and we deduce with the Cramer-Wold device, the continuous mapping theorem and Slutsky’s theorem that the second component of $\nabla\varphi_{n}$ multiplied by $\sqrt{n}$ also in distribution to a centered Gaussian random vector. It is immediate to deduce that the same convergence result holds for the third component, which is to say that $\sqrt{n}\left(\sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\right)$ converges in distribution to a centered Gaussian random variable. We finally deduce, with the Cramer-Wold device, that (iii) is fulfilled.
@@
$\mathbf{H}(\boldsymbol{\beta}_{-0})=$ $2\sum_{k=1}^{K}\pi_{k}\left[\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}+\frac{1}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}\right],$ (B.6)
which is a definite positive matrix under hypothesis $\mathbf{H}_{1}$. Using a block matrix determinant formula, we have
$\left|\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})\right|$ $=\begin{vmatrix}2&0\\ 0&\frac{1}{2}\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\end{vmatrix}\ \begin{vmatrix}\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}\end{vmatrix}$ (B.7)
where $\mathbf{C}=\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}&\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\end{pmatrix}$ and it only has to be verified that the second determinant at the righthand side of (B.7) is strictly positive. We now have to show that
@@
$+2\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}-\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}$ (B.8)
@@
using the fact that $\sum_{k}(\sqrt{\pi_{k}})^{2}=1$. It can be noted that if $\mathbf{u}\neq 0$, previous inequality is strict unless $\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{1}=\cdots=\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{K}$, which can not happen under hypothesis $\mathbf{H}_{1}$. The second part at the righthand side of (B.8) is handled the same way. We have
@@
This is a direct consequence of the continuous mapping theorem, which gives us that for all $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\mathcal{N}$, $\|\nabla_{00}\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\mathbf{H}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\|\to 0$ in probability, and the fact that third order partial derivatives of $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ are bounded in probability for $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ so that Theorem A.2 can apply. ∎
Proof.
of Lemma 5.1
@@
$=\frac{\mbox{Cov}_{w}(X,Y)}{\mbox{Var}_{w}(X)},$ (B.9)
with the notations $\mu_{X,w}=\sum_{k=1}^{K}w_{k}\mu_{X}^{k}$, $\mu_{Y,w}=\sum_{j=1}^{K}w_{j}\mu_{Y}^{j}$, $\mbox{Cov}_{w}(X,Y)=\sum_{k=1}^{K}w_{k}\mu_{X}^{k}\mu_{Y}^{k}-\mu_{X,w}\mu_{Y,w}$ and $\mbox{Var}_{w}(X)=\sum_{k=1}^{K}w_{k}(\mu_{X}^{k})^{2}-(\mu_{X,w})^{2}$. The gradient $\nabla g$ of $g$, evaluated at point $(\mu_{X}^{1},\ldots,\mu_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$, is equal to
@@
$=\frac{1}{(\mbox{Var}_{w}(X))^{2}}\sum_{k=1}^{K}w_{k}^{2}\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}\left(\beta_{1}^{2}\sigma^{2}_{X,k}+\sigma^{2}_{Y,k}\right)$ (B.10)

### v4

@@ 9 Discussion
This work focuses on a statistical framework designed to extract dependencies from experiments, specifically introducing linear regression estimators in the context where the predictor and predicted variables are not jointly observed but share a common observed categorical variable. In this work we have chosen the basic linear multivariate setting, prioritizing simplicity and computational feasibility. In particular, the estimator based on the method of moments makes no hypotheses about the data distribution and can be computed explicitly. The optimal transport estimator involves a simple optimization problem and is based on the Gaussian form of the Wasserstein distance, but does not technically require the data to be Gaussian, seeking to approximate them with Gaussian variables in any case. The proposed bootstrap procedure produces confidence intervals for the regression parameters that are smaller than those obtained with the naive approach, while preserving a high coverage rate. In practice, this allows better detection of significant effects in cases where the sample size is small, which is often the case in in vivo experiments.
However, these approaches are not applicable in cases where the linear relationship hypothesis cannot be satisfied. For example, this is the case when predicting survival data with some continuous biomarkers, which is of particular interest in the research on the adverse effects of radiation. To be able to consider such scenarios, our model can be extended to a more general case, namely with a generalized linear model. The optimal transport estimator seems promising in this context, given the fact that the Wasserstein distance allows to compare probability distributions of different nature (e.g. continuous and discrete).
Another potential direction of research is to investigate alternative methods based on integrated likelihood and Bayesian approaches, which are likely to produce better results in many cases, but require the imposition of priors on distributions.
Finally, it would be of interest to work on improving the theoretical properties of our finite sample estimators, namely the correction of the negative bias that appears for both estimators. The latter is particularly important in the case of the optimal transport estimator, which can arise naturally with the Wasserstein distance (e.g. Manole et al., (2024)). Correcting this bias would considerably improve the estimator, making it competitive with the aforementioned approaches, which make numerous assumptions about the data.
@@ Acknowledgements
The authors gratefully acknowledge the funding from the European Union through the PO FEDER-FSE Bourgogne 2014/2020 programs as part of the ModBioCan2020 project, and from the Institut de Radioprotection et de Sûreté nucléaire as part of the ROSIRIS project. The authors would also like to express their gratitude to Dr. Agnès François, Dr. Annaïg Bertho, Dr. Morgane Dos Santos and Dr. Fabien Milliat for the experimental data. Finally, the authors warmly thank Dr. Agnès François for her help in the biological interpretation of the data, as well as Dr. Patrick Tardivel for his numerous remarks that helped to improve the presentation of the manuscript.
Supporting Information
The Python code is freely available at github.com/parsenteva/vivo_lm. The data that support the findings in this paper are available upon request from Dr. Agnès François (agnes.françois@irsn.fr).
@@ Appendix A Some classical theorems in asymptotic statistics
Under additional hypotheses, we also get the asymptotic normality of the sequence of estimators $\widehat{\theta}_{n}$ of $\theta_{0}$. We denote by $\nabla_{00}Q_{n}(\theta)$ the Hessian matrix of the functional $Q_{n}$ evaluated at $\theta$.
Theorem A.4
@@
Theorem A.5
(Central limit theorem for the bootstrap means)
@@ Appendix B Proofs
The law of large numbers gives that $\widehat{\boldsymbol{\Gamma}}_{X}^{2}\to\boldsymbol{\Gamma}_{X}^{2}$ and $\widehat{\sigma}_{Y}^{2}\to\sigma_{Y}^{2}$ in probability and we deduce, with another application of the continuous mapping theorem, that $\widehat{\sigma}_{Y}^{2}-\widehat{\boldsymbol{\beta}}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{2}\widehat{\boldsymbol{\beta}}\to\sigma_{Y}^{2}-\boldsymbol{\beta}^{\top}\boldsymbol{\Gamma}_{X}^{2}\boldsymbol{\beta}=\sigma_{\epsilon}^{2}$ in probability as $n_{\min}$ tends to infinity. $\Box$
Proofof Lemma 4.2
@@
$\left|\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}-\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}\right|$ $\leq C_{2,k}\left\|\widehat{\boldsymbol{\Gamma}}_{X}^{k}\right\|_{sp}\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|.$ (11)
@@
$\left|\sqrt{\boldsymbol{\alpha}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\alpha}_{-0}+\sigma_{\alpha}^{2}}-\sqrt{\boldsymbol{\gamma}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\gamma}_{-0}+\sigma_{\gamma}^{2}}\right|$ $\leq B_{n}^{k}\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)$ (12)
@@
$\left|\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\varphi_{n}(\boldsymbol{\alpha},\sigma_{\alpha}^{2})\right|$ $\leq\left(\|\boldsymbol{\alpha}-\boldsymbol{\gamma}\|+|\sigma_{\alpha}^{2}-\sigma_{\gamma}^{2}|\right)\sum_{k=1}^{K}\pi_{k}\left(B_{n}^{k}+A_{n}^{k}\right),$ (13)
@@
We conclude the proof by recalling that $\varphi(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ attains its unique minimum at $(\boldsymbol{\beta},\sigma^{2}_{\epsilon})\in\Theta$ if assumption $\mathbf{H}_{1}$ is fulfilled, so that $(\widehat{\boldsymbol{\beta}}^{W},\widehat{\sigma}^{2,W})\to(\boldsymbol{\beta},\sigma^{2}_{\epsilon})$ in probability in view of Theorem 2.1 in Newey and McFadden, (1994). $\Box$
Proofof Proposition 4.1
@@
$\sqrt{n}\begin{pmatrix}\widehat{\boldsymbol{\mu}}_{X}^{1}-\boldsymbol{\mu}_{X}^{1}\\ \vdots\\ \widehat{\boldsymbol{\mu}}_{X}^{K}-\boldsymbol{\mu}_{X}^{K}\\ \widehat{\mu}_{Y}^{1}-\mu_{Y}^{1}\\ \vdots\\ \widehat{\mu}_{Y}^{K}-\mu_{Y}^{K}\end{pmatrix}$ $\rightsquigarrow\mathcal{N}\left(0,\boldsymbol{\Gamma}_{\mu}\right)$ (14)
@@
Application $g$ is differentiable at $\boldsymbol{\theta}=(\boldsymbol{\mu}_{X}^{1},\ldots,\boldsymbol{\mu}_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$, with non null Jacobian matrix denoted by $\mathbf{J}_{\theta}$ (see Chapter 8 and more particularly Theorem 8.3 in Magnus and Neudecker, (2019)). The application of the Delta method (see Theorem 3.1 in van der Vaart, (1998)) allows to get the asymptotic normality convergence result,
@@
where $\boldsymbol{\Gamma}_{\beta_{M}}=\mathbf{J}_{\theta}\boldsymbol{\Gamma}_{\mu}\mathbf{J}_{\theta}^{\top}$. $\Box$
Proofof Proposition 4.2
@@
The proof consists in checking the different points of Theorem A.4. Point (i) is satisfied by the assumptions, and the point (ii) follows directly from the fact that $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2})$ is twice-differentiable in a neighborhood of $(\boldsymbol{\beta},\sigma_{\epsilon}^{2})$. To show that (iii) is fulfilled, we consider the following expansion, based on the empirical version of the gradient of $\varphi$:
@@
$\nabla\varphi_{n}$ $=\begin{pmatrix}-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\\ -2\sum_{k=1}^{K}\pi_{k}\left[\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\widehat{\boldsymbol{\mu}}_{X}^{k}+\left(\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}-1\right)\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}\right]\\ \sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\end{pmatrix}$ (15)
@@
Since model (1) holds, $\nabla\varphi=0$ and $\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}=(\widehat{\mu}_{Y}^{k}-{\mu}_{Y}^{k})-\boldsymbol{\beta}_{-0}^{\top}\left(\widehat{\boldsymbol{\mu}}_{X}^{k}-\boldsymbol{\mu}_{X}^{k}\right)$, we thus deduce with (14) the asymptotic normality of the first component of the gradient $\nabla\varphi_{n}$, that is to say $\sqrt{n}\left(-2\sum_{k=1}^{K}\pi_{k}\left(\widehat{\mu}_{Y}^{k}-\beta_{0}-\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\mu}}_{X}^{k}\right)\right)$ converges in distribution to a centered Gaussian distribution. As far as the second component is concerned, it can be noted that $\widehat{\boldsymbol{\Gamma}}_{X}^{k}$ converges in probability to $\boldsymbol{\Gamma}_{X}^{k}$, and by the continuous mapping theorem $\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}\to\sigma_{Y,k}$ in probability. It can also be noted that, under the moment condition ${\mathbb{E}}\left[\|\mathbf{X}\|^{4}|G=k\right]<\infty$, the central limit theorem gives that $\sqrt{n}\left(\widehat{\boldsymbol{\Gamma}}_{X}^{k}-\boldsymbol{\Gamma}_{X}^{k}\right)$ converges in distribution to a centered Gaussian multivariate distribution, and we deduce with the Cramer-Wold device, the continuous mapping theorem and Slutsky’s theorem that the second component of $\nabla\varphi_{n}$ multiplied by $\sqrt{n}$ also in distribution to a centered Gaussian random vector. It is immediate to deduce that the same convergence result holds for the third component, which is to say that $\sqrt{n}\left(\sum_{k=1}^{K}\pi_{k}\left(1-\frac{\widehat{\sigma}_{Y,k}}{\sqrt{\boldsymbol{\beta}_{-0}^{\top}\widehat{\boldsymbol{\Gamma}}_{X}^{k}\boldsymbol{\beta}_{-0}+\sigma_{\epsilon}^{2}}}\right)\right)$ converges in distribution to a centered Gaussian random variable. We finally deduce, with the Cramer-Wold device, that (iii) is fulfilled.
@@
$\mathbf{H}(\boldsymbol{\beta}_{-0})=$ $2\sum_{k=1}^{K}\pi_{k}\left[\boldsymbol{\mu}_{X}^{k}\left(\boldsymbol{\mu}_{X}^{k}\right)^{\top}+\frac{1}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}\right],$ (16)
which is a positive definite matrix under the hypothesis $\mathbf{H}_{1}$. Using a block matrix determinant formula, we have
$\left|\mathbf{H}(\boldsymbol{\beta},\sigma_{\epsilon}^{2})\right|$ $=\begin{vmatrix}2&0\\ 0&\frac{1}{2}\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\end{vmatrix}\ \begin{vmatrix}\mathbf{H}(\boldsymbol{\beta}_{-0})-\mathbf{C}\begin{pmatrix}\frac{1}{2}&0\\ 0&\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\end{pmatrix}\mathbf{C}^{\top}\end{vmatrix}$ (17)
where $\mathbf{C}=\begin{pmatrix}2\sum_{k=1}^{K}\pi_{k}\boldsymbol{\mu}_{X}^{k}&\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\end{pmatrix}$, and it only has to be verified that the second determinant at the righthand side of (17) is strictly positive. We now have to show that
@@
$+2\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\left(\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}-\frac{2}{\sum_{k}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}}\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)\left(\sum_{k=1}^{K}\frac{\pi_{k}}{\sigma_{Y,k}^{2}}\boldsymbol{\Gamma}_{X}^{k}\boldsymbol{\beta}_{-0}\right)^{\top}$ (18)
@@
using the fact that $\sum_{k}(\sqrt{\pi_{k}})^{2}=1$. It can be noted that if $\mathbf{u}\neq 0$, previous inequality is strict unless $\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{1}=\cdots=\mathbf{u}^{\top}\boldsymbol{\mu}_{X}^{K}$, which cannot happen under the hypothesis $\mathbf{H}_{1}$. The second part at the righthand side of (18) is handled in the same way. We have
@@
This is a direct consequence of the continuous mapping theorem, which gives us that for all $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\in\mathcal{N}$, $\|\nabla_{00}\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})-\mathbf{H}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})\|\to 0$ in probability, and the fact that third order partial derivatives of $\varphi_{n}(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ are bounded in probability for $(\boldsymbol{\gamma},\sigma^{2}_{\gamma})$ so that Theorem A.2 can apply. $\Box$
Proofof Lemma 5.1
@@
$=\frac{\mbox{Cov}_{w}(X,Y)}{\mbox{Var}_{w}(X)},$ (19)
with the notations $\mu_{X,w}=\sum_{k=1}^{K}w_{k}\mu_{X}^{k}$, $\mu_{Y,w}=\sum_{j=1}^{K}w_{j}\mu_{Y}^{j}$, $\mbox{Cov}_{w}(X,Y)=\sum_{k=1}^{K}w_{k}\mu_{X}^{k}\mu_{Y}^{k}-\mu_{X,w}\mu_{Y,w}$ and $\mbox{Var}_{w}(X)=\sum_{k=1}^{K}w_{k}(\mu_{X}^{k})^{2}-(\mu_{X,w})^{2}$. The gradient $\nabla g$ of $g$, evaluated at the point $(\mu_{X}^{1},\ldots,\mu_{X}^{K},\mu_{Y}^{1},\ldots,\mu_{Y}^{K})$, is equal to
@@
$=\frac{1}{(\mbox{Var}_{w}(X))^{2}}\sum_{k=1}^{K}w_{k}^{2}\left(\mu_{X}^{k}-\mu_{X,w}\right)^{2}\left(\beta_{1}^{2}\sigma^{2}_{X,k}+\sigma^{2}_{Y,k}\right)$ (20)